On a rough day there is a uniform current in the river,the speed of da river is v .which will help the onward journey and impede the return.if the boat takes t sec t" go and come back .wat is ratio between t"/t

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear akhil

question is not clear

let speed of boat is u

t=d/(v+u)

t'' =d/(u-v)

so t/t''  =(u-v)/(u+v)

Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin